After 6 games played in a competition, Mavericks Football Club averaged 1.5 goals per game. How many goals would they have to score in the seventh game for that average to increase to 2 goals per game? _______



1
Expert's answer
2021-12-15T17:18:56-0500

If it is 1.5 average for 6 observation, then totally 6*1.5 = 9 goals was scored. To make an average of 2 goals in 7 games the team should score 2*7 = 14 times in 7 games, which means they need to score 14-9=5 goals in the last(seventh) match